OpenStudy (anonymous):

if x=17 find 1/9 of x 1/3 of x 1/2 of x answer cant be in decimal.. and show how you did it.

OpenStudy (anonymous):

17 is prime so none of the fractions can be cancelled down. So the answers are: 17/9, 17/3 and 17/2 respectively.

OpenStudy (anonymous):

if i say that there are 17 elephants.. and it is to be distributed among 3 people in the said ratio..i.e. 1/9,1/3,1/2. so obviously elephants cant be in decimals.but this problem has a solution think laterally

OpenStudy (anonymous):

1/9 of x if x=17 1/9 x 17= 17/9 1/3 of x 1/3 x 17 17/3 1/2 x 17= 17/2

OpenStudy (anonymous):

oh i know this add 1 elephant so that the total is 18 then subtract 1/9 which is 2 1/3 is 6 1/2 is 9 you are left with 18 - 9 =9 -6=3 - 2 = 1 that is the elephant you added

OpenStudy (anonymous):

shrishti..suppose you have to divide 17 elephants in the given ratio..canthere be 17/9 elephants

OpenStudy (anonymous):

is that right ?

OpenStudy (anonymous):

good ishan

OpenStudy (anonymous):

Avadh , from where elephant came here?

OpenStudy (anonymous):

hehe..in the original question posted..i wrote that answer cannot be in decimal.so i just gave u an example to clarify that..

OpenStudy (anonymous):

seriously mujhe tou tum samajh nahi aate aur nahi tumhare question.

OpenStudy (anonymous):

please take a look at my questions :)